# GTPC-SSCD: Gate-guided Two-level Perturbation Consistency-based Semi-Supervised Change Detection thanks: This work was supported in part by the Scientific Research Program of Tianjin Municipal Education Commission under Grant 2023KJ232. (Corresponding author: Rui Huang.)

### v1

@@ III-A Experimental Setups
WHU-CD 1 $15354\times 32507\times 3$ $5974/743/744$ 0.2 m/pixel building
LEVIR-CD 637 $1024\times 1024\times 3$ $7120/1024/2048$ 0.5 m/pixel building
@@
GZ-CD 19 $1006\times 1168\times 3,$ $4936\times 5224\times 3$ $2882/360/361$ 0.55 m/pixel building
@@
Criterion. Following Bandara et al. [15], Mao et al. [19], and Yang et al. [23], we adopt intersection over union (IoU) and overall accuracy (OA) as the primary performance metric to evaluate different change detection methods.
TABLE II: Quantitative comparison of different methods on six CD datasets. The highest scores are marked in bold.
@@
Figure 3: Detection results of different methods on six CD datasets at the $5\%$ labeled training ratio. Different colors are used for a better view, i.e., white for true positive, black for true negative, red for false positive, and green for false negative.
@@ III-B Results and Discussion
Comparison to the State-of-the-Art. Table II exhibits the quantitative results of different methods on WHU-CD, LEVIR-CD, BCD, GZ-CD, EGY-BCD, and CDD datasets. From the findings in these tables, we can draw the following observations. ❶ Most SSCD methods outperform Sup-only under the same partition, confirming the benefit of unlabeled data for SSCD. ❷ Our method achieves the best performance across all four partitions, notably surpassing other methods, especially on WHU-CD and GZ-CD. On WHU-CD, compared to the current SOTA method UniMatch, our method brings $4.3\%$, $3.2\%$, $3.0\%$, and $1.7\%$ performance gain in terms of IoU with $5\%$, $10\%$, $20\%$, and $40\%$ labeled training data, respectively. On GZ-CD, the improved performance with IoU of our method over the best UniMatch are $5.2\%$, $1.2\%$, $5.0\%$, and $7.3\%$.
Detection results. Fig. 3 shows some examples of different methods on six CD datasets under the partition of $5\%$. Our method exhibits higher accuracy and richer details. Both quantitative and qualitative results support our method’s superiority.
Computational complexity analysis. Table III shows a detailed comparison of parameters and computation costs of different methods. Compared to other methods, our method achieves a high FPS with a moderate parameter count, striking an effective balance between performance and efficiency.
TABLE III: Comparison of parameters and computational costs of different methods.
@@
Sensitivity of the gate mechanism. We apply the optimal threshold on WHU-CD due to space limitations. The experimental results are shown in Fig 4. High ratio lead to unstable training as more samples are perturbed, while low ratio may miss perturbation opportunities. Median threshold strikes a balance.
Figure 4: Performance comparison with different perturbation ratios under 5% labeled training data on WHU-CD dataset.
@@ IV Conclusion
In this work, we introduce a novel gate-guided two-level perturbation consistency regularization-based SSCD method. This approach ensures image-level strong-to-weak consistency and feature-level perturbation consistency, enabling efficient utilization of a large amount of unlabeled data. Moreover, we design a gate module to intelligently access the need of feature perturbations. Extensive experiments on six benchmark datasets validate the effectiveness and superiority of our proposed method. In the future, we will further explore enabling the network to autonomously determine the most effective perturbation strategies tailored to various data scenarios.

### v2

@@ III-A Experimental Setup
WHU-CD 1 $15354\times 32507\times 3$ $5974/743/744$ 0.075 m/pixel building
LEVIR-CD 637 $1024\times 1024\times 3$ $7120/1024/2048$ 0.5 m/pixel building, villa residence, apartment, garage, etc.
@@
GZ-CD 19 $1006\times 1168\times 3,$ $4936\times 5224\times 3$ $2882/360/361$ 0.55 m/pixel building, road, forests, farmlands, etc.
@@
TABLE II: Quantitative comparison of different methods on six CD datasets. The highest scores are in bold, and the second are underlined.
@@
Datasets. We conduct experiments on six widely used remote sensing CD datasets, namely WHU-CD [18], LEVIR-CD [25], BCD [18], GZ-CD [6], EGY-BCD [26], and CDD [27]. Table I shows the details of these datasets, including the number of image pairs, image size, train/val/test splits, spatial resolution, and change types. GZ-CD and CDD datasets encompass a diverse range of change types, such as buildings, vehicles, and roads, and others. The remaining four CD datasets mainly concentrate on various forms of building alterations. All images are cropped into non-overlapping patches of size $256\times 256$, which are then divided into training, validation, and test sets. The training set is further divided into labeled and unlabeled data with the following ratios: $[5\%,95\%]$, $[10\%,90\%]$, $[20\%,80\%]$, $[40\%,60\%]$.
Criterion. Following Bandara et al. [24] and Yang et al. [17], we adopt intersection over union (IoU) and overall accuracy (OA) as evaluation metrics. IoU measures the overlap between predicted and actual changes for precision, and OA gives an overall accuracy measure. Their formulas are:
$\mathrm{IoU}=\mathrm{TP}/(\mathrm{TP+FP+FN})$ (14)
$\mathrm{OA}=(\mathrm{TP+TN})/(\mathrm{TP+FP+FN+TN})$ (15)
where TP, TN, FP, and FN represent true positive, true negative, false positive, and false negative, respectively.
Figure 3: Detection results of different methods on six CD datasets at $5\%$ labeled ratio. Here, white for TP, black for TN, red for FP, and green for FN.
@@ III-B Results and Discussion
Comparison with the state-of-the-art. Table II exhibits the quantitative results of different methods on WHU-CD, LEVIR-CD, BCD, GZ-CD, EGY-BCD, and CDD datasets. From the findings in these tables, we can draw the following observations. ❶ Most SSCD methods outperform Sup-only under the same partition, confirming the benefit of unlabeled data for SSCD. ❷ Our method achieves the best performance across all four partitions, notably surpassing other methods, especially on WHU-CD and GZ-CD. On WHU-CD, compared to the leading method UniMatch, our method brings $4.3\%$, $3.2\%$, $3.0\%$, and $1.7\%$ performance gain in terms of IoU with $5\%$, $10\%$, $20\%$, and $40\%$ labeled training data, respectively. On GZ-CD, the improved performance with IoU of our method over the best UniMatch are $5.2\%$, $1.2\%$, $5.0\%$, and $7.3\%$.
Detection results. Fig. 3 shows some examples of different methods on six CD datasets under $5\%$ labeled partition. Our method exhibits higher accuracy and richer details. Both quantitative and qualitative results support our method’s superiority.
Computational complexity analysis. Table III details the parameters and computation costs of different methods. Our method shows 65.87 GFLOPs, 57.3 M parameters, and an FPS of 51.64. Compared with UniMatch, our method significantly boosts performance with moderate GFLOPs and parameters, balancing performance and efficiency effectively.
TABLE III: Comparison of parameters and computational costs.
@@
Figure 4: Comparison of different perturbation ratios at $5\%$ labeled WHU-CD.
Sensitivity of the gating mechanism. Fig. 4 shows the perturbation ratio experiments on WHU-CD with $5\%$ labeled data. A high ratio leads to unstable training as more samples are perturbed, while a low ratio may miss perturbation opportunities. The median ratio achieves a proper balance and attains the best performance.
@@ IV Conclusion
In this work, we introduce the GTPC-SSCD, a novel semi-supervised change detection method based on gate-guided two-level perturbation consistency regularization. It maintains both image-level strong-to-weak consistency and feature-level perturbation consistency, facilitating efficient utilization of the abundant unlabeled data. We also develop a hardness analysis-based gating mechanism to intelligently determine the necessity of implementing feature perturbations per sample. Extensive experiments on six benchmark change detection datasets validate the effectiveness and superiority of our proposed method. In the future, we plan to explore network autonomy in perturbation strategy selection for diverse data scenarios to boost the performance and applicability.